OpenStudy (anonymous):

What is a Parabola with vertex at (0, 0) and directrix at x = 5 look like written in standard form?

OpenStudy (anonymous):

Give standard form.

OpenStudy (goformit100):

@Itz_U_Girl

terenzreignz (terenzreignz):

Standard form is (x - h)^2 = 4p(y - k), where (h,k) is the vertex, and p is the perpendicular distance of said vertex from the directrix. the vertex being (0,0) makes things easy, it becomes x^2 = 4py, The distance of the vertex (0,0) from the directrix is 5, so |p|=5, but Since the directrix is above the vertex, that suggests the parabola is concave downward, so.. p = -5 x^2 = -20y (not entirely sure about this, your comments/corrections are welcome)